the height of the flagpole is three fourths the height of the school. the difference in there heights is 4.5 m. what is the height of the school?
Answer by stanbon(75887)

the height of the flagpole is three fourths the height of the school. the difference in there heights is 4.5 m. what is the height of the school?
----------
Equation:
f = (3/4)s
s-f = 4.5
----
Substitute for "f" and solve for "s":
s-(3/4)s = 4.5
(1/4)s = 4.5
s = 4*4.5
s = 18 meters (height of the school)
---------------
Since f = (3/4)s, f = (3/4)(18)
= (3/2)(9)= 13.5 meters (height of the flag pole)
===============================
Cheers,
Stan H.
=============